A recent survey found that

87​%

of​ first-year college students were involved in volunteer work at least occasionally. Suppose a random sample of

11

college students is taken. Find the probability that exactly

7

students volunteered at least occasionally.
Answer #1

þ » probability of student volunteered q o probability of student not volunteered nell, 8=7, p = (0-87) , 2= (0-13) P(x=8) =